# This draft: December 17, 2025 First draft: March 5, 2025

### v1

@@
This draft: December 17, 2025
@@ A Consensus-Bottleneck Asset Pricing Model
We introduce the Consensus-Bottleneck Asset Pricing Model (CB-APM), a partially interpretable neural network that replicates the reasoning processes of sell-side analysts by capturing how dispersed investor beliefs are compressed into asset prices through a consensus formation process. By modeling this “bottleneck” to summarize firm- and macro-level information, CB-APM not only predicts future risk premiums of U.S. equities but also links belief aggregation to expected returns in a structurally interpretable manner. The model improves long-horizon return forecasts and outperforms standard deep learning approaches in both predictive accuracy and explanatory power. Comprehensive portfolio analyses show that CB-APM’s out-of-sample predictions translate into economically meaningful payoffs, with monotonic return differentials and stable long-short performance across regularization settings. Empirically, CB-APM leverages consensus as a regularizer to amplify long-horizon predictability and yields interpretable consensus-based components that clarify how information is priced in returns. Moreover, regression and GRS-based pricing diagnostics reveal that the learned consensus representations capture priced variation only partially spanned by traditional factor models, demonstrating that CB-APM uncovers belief-driven structure in expected returns beyond the canonical factor space. Overall, CB-APM provides an interpretable and empirically grounded framework for understanding belief-driven return dynamics.
@@ 1 Introduction
Recent studies in machine learning asset pricing frequently employ models that are not interpretable, which raises concerns about relying on complex machine learning algorithms in empirical asset pricing without a clear understanding of why and how these models arrive at their conclusions. Furthermore, these papers often attempt to interpret the prediction results based on the learned models and derive economic implications. However, rudin2019stop argues that such analyses are solely based on post-hoc explanations that should be considered as fitting narratives to the outcomes. These explanations are conveniently aligned with prevailing economic theories and tend to disregard contradictory evidence, which limit the scope and applicability of the findings. For these reasons, rudin2019stop strives to rectify researchers and practitioners to use interpretable models over black-box algorithms. Despite growing interest in interpretable machine learning and trustworthy Artificial Intelligence (AI), a notable gap persists in applying and validating these approaches within asset pricing, beyond traditional regression or decision-tree models. In particular, existing machine-learning frameworks rarely achieve both strong predictive performance and economic interpretability. To address this gap, we propose the Consensus-Bottleneck Asset Pricing Model (CB-APM), a framework that employs a partially interpretable neural architecture to predict future stock returns while preserving clear economic structure.
@@
The remainder of the paper is organized as follows. Section 2 reviews interpretable machine learning and situates CB-APM within this framework. Section 3 outlines the model, estimation procedure, and architecture. Section 4 describes the data and evaluation design, including the autoencoder for macroeconomic state extraction. Section 5 presents empirical results on predictive performance, macroeconomic embeddings, and portfolio-based pricing implications. Section 6 investigates the pricing content of the approximated consensuses using regression and GRS tests. Section 7 concludes. The Internet Appendix provides additional robustness analyses and supplementary results.
2 Interpretable Artificial Intelligence
Nevertheless the black box nature of deep neural networks, researchers may attempt to make such models understandable by introducing the concept of “disentangled representations” (See higgins2018towards; locatello2019challenging; locatello2019fairness). To understand the concept of disentanglement, we can consider the simple example of a naïve feedforward neural network, which consists of multiple hidden layers. The hidden layers are intermediate vector representations, often interpreted as features extracted from input data. Since they are trained with a downstream task-related objective in most cases, those layers are presumed to represent the essential information within the high-dimensional and noisy input data, well enough for the successful performance on that task. However, these features exist in an “entangled” manner, meaning that each element of the feature representation is a mixture of multiple factors. This entanglement complicates the interpretability of the model, as it blurs the specific contributions of individual input features to the final output. In this context, “disentanglement” refers to separating the underlying causal factors of the data and intermediate representations into distinct and non-overlapping representations. For instance, in fixed income markets, dozens of yields across maturities can be effectively summarized by three disentangled factors, level, slope, and curvature, as illustrated in Figure 1. Each of these dimensions captures an independent source of variation in the yield curve, and together they provide an interpretable low-dimensional representation that still preserves the essential structure of the data. [^2] A clarification is needed to avoid conflating our framework with traditional factor models. nelson1987parsimonious impose level, slope, and curvature terms to parameterize the yield curve, while litterman1991common show via principal components that similar dimensions emerge empirically. Both approaches reduce high-dimensional yields to a few interpretable coordinates. By contrast, in CB-APM the consensus-bottleneck is not imposed ex ante but learned endogenously, closer in spirit to Litterman–Scheinkman than to Nelson–Siegel. In this sense, whereas Nelson–Siegel estimate factors to fit the curve, our model embeds interpretability within the network architecture itself, rendering part of the neural network transparent while retaining predictive capacity.
[Insert Figure 1 here]
@@ 3.1 Return prediction model
$R_{i,t+h}=E_{t}[R_{i,t+h}]+\epsilon_{i,t+h},$ (1)
where $R_{i,t+h}$ is the $h$-month return of asset $i$ excess of the risk-free rate at time $t+h$, and $\epsilon_{i,t+h}$ is an error term. In this context, $h$ is used to assign the forecasting horizon, enabling the consideration of multi-horizon predictions, allowing CB-APM to model long-term dependencies. The expected excess return in equation (1) is defined as the expectation conditional on information sets,
$E_{t}[R_{i,t+h}]=E[R_{i,t+h}|I^{f}_{i,t},I^{m}_{t}].$
Here, $I^{f}_{i,t}$ and $I^{m}_{t}$ are the set of firm-specific characteristics and macroeconomic predictors at time $t$, respectively.  [^4] A detailed description of the predictors comprising the information sets is provided in Section 4, and a complete list of variables is available in Internet Appendix D. The macroeconomic information set $I^{m}_{t}$ is represented empirically by a latent vector extracted through an autoencoder trained on macroeconomic variables, as described in Section 4.2. It is important to note that consensus information is deliberately excluded from $I^{f}_{i,t}$.
@@
$E[R_{i,t+h}|I^{f}_{i,t},I^{m}_{t}]=g(f(I^{f}_{i,t},I^{m}_{t};\phi);\theta),$ (2)
where the function $f(\cdot)$ and $g(\cdot)$ are smooth functions parameterized by learnable parameters $\theta$ and $\phi$. The function $f(\cdot)$ is specifically designed to model the conditional expectation of analyst consensus. Then, the function $g(\cdot)$ models the expected return only using the features of approximated consensus from the previous step, creating a “concept-bottleneck” within the prediction model. This empirical design is predicated on the understanding that both researchers in empirical asset pricing and financial analysts share the objective of assessing a firm’s value and discerning the factors that influence these valuations. While analysts often have access to broader datasets, including some predictive signals that may not be publicly available or included in this article, asset pricing panel data can represent a information subset in a descent quality by providing a comprehensive and quantifiable measures of firm’s fundamentals and macroeconomic conditions that are crucial for the approximation of the consensus, as shown in the empirical results later on.
@@
$E_{t}[R_{i,t+h}]=g(\hat{C}_{i,t};\theta).$ (3)
@@
Although it is designed as intended, given that neural networks are well-recognized as “universal approximators”, the model can allow any scenarios and consequences as outcomes, that doesn’t necessarily align with the economic theories. To overcome the limitation of the proposed prediction model, we apply stabilized optimization approaches proposed in the machine learning literature, such as regularization and scheduling. Such techniques are expected to function as “universal constraints”, achieving both practical performances and theoretical rigor. See Internet Appendix B.3 for detailed discussions and experimental settings.
@@ 3.2 Estimation
$L=L_{R}+\lambda L_{C},$ (4)
where the “return loss” $L_{R}$ is formulated as,
$L_{R}(\phi,\theta)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(R_{i,t+h}-g(f(I^{f}_{i,t},I^{m}_{t};\phi);\theta))^{2},$ (5)
and the “consensus loss” $L_{C}$ is formulated as,
$L_{C}(\phi)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(C_{i,t}-f(I^{f}_{i,t},I^{m}_{t};\phi))^{2}.$ (6)
$L_{R}$ and $L_{C}$ are cross-sectional mean squared errors (MSE) from a standard pooled OLS estimator, where $\lambda$ is a hyperparameter that assigns weight to the consensus loss, providing additional flexibility into the empirical design of the model.
@@
The strategy of jointly learning the consensus and excess return offers several advantages. Firstly, it tends to yield higher performance metrics due to the synergistic learning of interconnected variables. An alternative method might involve independent optimization, where the $f(\cdot)$ and $g(\cdot)$ are trained independently in two separate steps. However, this segmented approach often fails to capture the potential inter-dependencies between the consensus estimates and the resulting excess returns. Furthermore, since the information set $I^{f}_{i,t}$ and $I^{m}_{t}$ are not included in equation (5), the training of $g(\cdot)$ entirely depend on the quality of the extracted signals in approximated consensus, which makes training with the loss function $L_{R}$ extremely challenging.
Secondly, it provides deeper insights and more intuitive understanding of the underlying financial dynamics. Independently learning $f(\cdot)$ using the equation (6) is not a novel concept and aligns with the existing literature supporting the evidence of the rational expectations hypothesis. As discussed in previous sections, the set of predictor signals used in this study is regarded to contain a significant amount of information sufficient to make “rational” expectations, [^5] Appendix C.5.2 formally evaluates this claim by examining the consensus-only specification corresponding to $\lambda\to\infty$ in Equation (4), showing that the model learns analysts’ consensus variables remarkably well (out-of-sample $R^{2}=30.30\%$) even without any return-prediction objective. This validates the architectural design of the consensus-bottleneck and provides empirical support for the rational expectations interpretation underlying the model. which simplifies the problem of approximating the opinions of individuals, compared to predicting future returns of assets. However, since analysts perform their analysis as their job, they must think and act beyond being merely “rational”; they must be “professional”. Therefore, we posit that professional and successful analysts strive to make their estimates predict not only “macroeconomic” consequences but also “firm-specific” outcomes. More specifically, proficient analysts will make decisions that better predict the future returns of a firm’s stocks.
@@ 3.3 Model architecture
In the proposed model, the consensus module is designed as an arbitrary feedforward network, while the prediction module is restricted to a simple linear regression that receives consensus variables as inputs and yield the expected excess return. This design choice is critical for enhancing interpretability in particular. When both modules are complex feedforward networks with multiple hidden layers, the advantage of using a consensus-based approach diminishes since it creates two separate black-box models from a single black-box model.
@@
$GELU(x)=x\cdot P(X\leq x)=x\cdot\Phi(x),$
where $\Phi(x)$ the cumulative distribution function for Gaussian distribution $X\sim N(0,\sigma^{2})$. GELU was first introduced by hendrycks2016gaussian as an alternative of Rectified linear units (ReLU) (nair2010rectified). Figure 3 shows that GELU permits some small interval for negative inputs to propagate through subsequent layers. See Internet Appendix B for a review of the literature on activation functions and justification for the selection of GELU.
@@
$H^{(1)}$ $=GELU(W_{0}(I^{f}\oplus I^{m})+b_{0})$
$H^{(2)}$ $=GELU(W_{1}H^{(1)}+b_{1})$
@@
$g(f(I^{f},I^{m};\phi);\theta)=W_{n}^{r}f(I^{f},I^{m};\phi)+b_{n}^{r}.$
@@ 4.1 Data description
In this section, we provide the brief explanations on the dataset and the sampling splitting scheme employed for empirical studies. The dataset comes from four distinct sources, which are all publicly available at the moment. Firstly, we obtain open-source asset pricing panel data from ChenZimmermann2021, available to download on their website (https://www.openassetpricing.com/). [^6] Data from ChenZimmermann2021 undergoes several preprocessing steps including lagging, data sampling, data imputation, and rank normalization, as detailed in Internet Appendix A. It comprises 114 firm-level predictors consisting of diverse financial metrics such as accounting figures, 13F filings, trading activities, and derivatives data.
@@
The final merged dataset consist of samples spanning from January 1994 to December 2023, with total 605,722 samples from 4,683 U.S. companies. Detailed descriptions of the dataset components and their respective sources are provided in Internet Appendix D.
@@ 4.2 Extracting macroeconomic state variables via autoencoder
To incorporate macroeconomic dynamics into the conditional expectation function $E_{t}[R_{i,t+h}]$ defined in equation (2), we encode the aggregate information set $I_{t}^{m}$ through an autoencoder-based representation. While macroeconomic variables are often dismissed in cross-sectional asset pricing due to their perceived homogeneity across firms, we argue that macro context exerts differentiated influence through sectoral dynamics, capital structure sensitivity, and behavioral channeling. However, the sheer volume and redundancy of macroeconomic indicators, particularly those sourced from databases such as FRED-MD, pose significant challenges for model training. Including hundreds of highly correlated variables not only increases the risk of overfitting but also dilutes the learning signal by overwhelming the model with noise and irrelevant information.
@@
Dimensionality reduction techniques have long been used in financial modeling to address such issues. Principal Component Analysis (PCA) has served as a standard tool for extracting latent factors from large panels of macroeconomic variables (ludvigson2007empirical), while extensions such as Sparse PCA and Independent Component Analysis (ICA) have been applied to improve factor interpretability and reduce multicollinearity (fan2016projected; erichson2020sparse). More recently, deep learning approaches—particularly autoencoders—have gained traction in the asset pricing literature for their ability to capture nonlinear interactions and extract economically meaningful latent structures from noisy, high-dimensional data (chen2024deep; gu2021autoencoder). These methods have proven effective in modeling complex macro-financial dynamics that traditional linear techniques may fail to uncover. [^8] Appendix C.4.2 demonstrates that replacing the autoencoder with a 32-factor PCA markedly weakens out-of-sample return predictability, despite both approaches delivering similar accuracy in reconstructing analysts’ consensus. This divergence highlights the advantage of nonlinear compression in capturing macroeconomic structure relevant for pricing.
@@
Formally, let the macroeconomic input at time $t$ be $\mathbf{x}_{t}\in\mathbb{R}^{D}$, where $D=123$. The encoder $f_{\phi}(\cdot)$ maps this input to a latent representation $\mathbf{z}_{t}\in\mathbb{R}^{d}$: [^9] Empirically, setting the latent dimension to $d=32$ yields the best out-of-sample performance (see Internet Appendix C.4.1).
$\mathbf{z}_{t}=f_{\phi}(\mathbf{x}_{t}).$
The decoder $g_{\theta}(\cdot)$ reconstructs the input:
$\hat{\mathbf{x}}_{t}=g_{\theta}(\mathbf{z}_{t}),$
@@
$L_{\text{AE}}(\theta,\phi)=\frac{1}{T}\sum_{t=1}^{T}\left\|\mathbf{x}_{t}-g_{\theta}(f_{\phi}(\mathbf{x}_{t}))\right\|_{2}^{2}.$
@@
Formally, the latent representation $\mathbf{z}_{t}$ learned through the autoencoder serves as an empirical proxy for the macroeconomic information set $I_{t}^{m}$ introduced in equation (2). In this context, $\mathbf{z}_{t}$ functions as a compressed, data-driven approximation of the macroeconomic state observable to investors at time $t$. This design allows the CB-APM to integrate the high-dimensional macroeconomic information set into a tractable latent representation, ensuring that firm-level forecasts remain conditioned on a parsimonious yet informative depiction of the aggregate economic environment. By mapping $I_{t}^{m}$ into $\mathbf{z}_{t}$, the model effectively operationalizes the theoretical information set within a learnable structure, thereby linking the empirical implementation of the macro encoder to the conditional expectation framework defined in equation (2).
As illustrated in Figure 4, this framework visualizes the overall data pipeline of the CB-APM, depicting how macroeconomic inputs are encoded, compressed, and subsequently integrated with firm-level characteristics for return prediction. The figure serves as a conceptual representation clarifying the interaction between the macro autoencoder and the return-prediction module. The full architecture details of the autoencoder, including hidden-layer configurations and activation functions, are provided in Internet Appendix B.3. The empirical findings underscore the importance of representing macroeconomic regimes in shaping cross-sectional return dynamics and highlight the utility of neural representation learning in extracting economically salient signals from high-dimensional macro data. At each expanding-window step, the autoencoder is trained only on macro data available up to the window end date, and the encoder is then used to compute $\mathbf{z}_{t}$ for that window’s validation and test months, thereby preventing look-ahead bias.

### v2

# This draft: December 23, 2025 First draft: March 5, 2025

@@
This draft: December 23, 2025
@@ A Consensus-Bottleneck Asset Pricing Model
We introduce the Consensus-Bottleneck Asset Pricing Model (CB-APM), a partially interpretable neural network that replicates the reasoning processes of sell-side analysts by capturing how dispersed investor beliefs are compressed into asset prices through a consensus formation process. By modeling this “bottleneck” to summarize firm- and macro-level information, CB-APM not only predicts future risk premiums of U.S. equities but also links belief aggregation to expected returns in a structurally interpretable manner. The model improves long-horizon return forecasts and outperforms standard deep learning approaches in both predictive accuracy and explanatory power. Comprehensive portfolio analyses show that CB-APM’s out-of-sample predictions translate into economically meaningful payoffs, with monotonic return differentials and stable long-short performance across regularization settings. Empirically, CB-APM leverages consensus as a regularizer to amplify long-horizon predictability and yields interpretable consensus-based components that clarify how information is priced in returns. Moreover, regression and Gibbons–Ross–Shanken (GRS)-based pricing diagnostics reveal that the learned consensus representations capture priced variation only partially spanned by traditional factor models, demonstrating that CB-APM uncovers belief-driven structure in expected returns beyond the canonical factor space. Overall, CB-APM provides an interpretable and empirically grounded framework for understanding belief-driven return dynamics.
@@ 1 Introduction
Recent studies in machine learning asset pricing frequently employ models that are not interpretable, [^2] While the finance literature often uses the terms “explainability” and “interpretability” interchangeably, we maintain a strict conceptual distinction to ensure the reliability of our economic inferences. To keep the focus on the economic mechanisms of belief aggregation, we provide a detailed taxonomy of AI transparency in Internet Appendix A. This section clarifies why our framework prioritizes interpretability-by-design over the post-hoc explainability (XAI) common in recent black-box models. Such rigor is essential to verify that the CB-APM is capturing priced fundamental signals rather than suffering from the “Clever Hans” phenomenon, where models achieve high accuracy through spurious, non-economic correlations. which raises concerns about relying on complex machine learning algorithms in empirical asset pricing without a clear understanding of why and how these models arrive at their conclusions. Furthermore, these papers often attempt to interpret the prediction results based on the learned models and derive economic implications. However, rudin2019stop argues that such analyses are solely based on post-hoc explanations that should be considered as fitting narratives to the outcomes. These explanations are conveniently aligned with prevailing economic theories and tend to disregard contradictory evidence, which limits the scope and applicability of the findings. For these reasons, rudin2019stop strives to rectify researchers and practitioners to use interpretable models over black-box algorithms. Despite growing interest in interpretable machine learning and trustworthy Artificial Intelligence (AI), a notable gap persists in applying and validating these approaches within asset pricing, beyond traditional regression or decision-tree models. In particular, existing machine-learning frameworks rarely achieve both strong predictive performance and economic interpretability. To address this gap, we propose the Consensus-Bottleneck Asset Pricing Model (CB-APM), a framework that employs a partially interpretable neural architecture to predict future stock returns while preserving clear economic structure.
@@
The remainder of the paper is organized as follows. Section 2 outlines the model, estimation procedure, and architecture. Section 3 describes the data and evaluation design, including the autoencoder for macroeconomic state extraction. Section 4 presents empirical results on predictive performance, macroeconomic embeddings, and portfolio-based pricing implications. Section 5 investigates the pricing content of the approximated consensuses using regression and GRS tests. Section 6 concludes. The Internet Appendix provides additional robustness analyses and supplementary results.
@@ 2.1 Return prediction model
$R_{i,t+h}=\mathbb{E}_{t}\!\left[R_{i,t+h}\right]+\varepsilon_{i,t+h},$ (1)
where $R_{i,t+h}$ is the $h$-month return of asset $i$ excess of the risk-free rate at time $t+h$, and $\varepsilon_{i,t+h}$ is an error term. In this context, $h$ is used to assign the forecasting horizon, enabling the consideration of multi-horizon predictions, allowing CB-APM to model long-term dependencies. The expected excess return in equation (1) is defined as the expectation conditional on information sets,
$\mathbb{E}_{t}\!\left[R_{i,t+h}\right]=\mathbb{E}\!\left[R_{i,t+h}\mid\mathcal{I}^{f}_{i,t},\,\mathcal{I}^{m}_{t}\right].$
Here, $\mathcal{I}^{f}_{i,t}$ and $\mathcal{I}^{m}_{t}$ are the sets of firm-specific characteristics and macroeconomic predictors at time $t$, respectively.  [^3] A detailed description of the predictors comprising the information sets is provided in Section 3, and a complete list of variables is available in Internet Appendix E. The macroeconomic information set $\mathcal{I}^{m}_{t}$ is represented empirically by a latent vector extracted through an autoencoder trained on macroeconomic variables, as described in Section 3.2. It is important to note that consensus information is deliberately excluded from $\mathcal{I}^{f}_{i,t}$.
@@
$\mathbb{E}\!\left[R_{i,t+h}\mid\mathcal{I}^{f}_{i,t},\,\mathcal{I}^{m}_{t}\right]=g\!\left(f\!\left(\mathcal{I}^{f}_{i,t},\,\mathcal{I}^{m}_{t};\phi\right);\theta\right),$ (2)
where the function $f(\cdot)$ and $g(\cdot)$ are smooth functions parameterized by learnable parameters $\theta$ and $\phi$. The function $f(\cdot)$ is specifically designed to model the conditional expectation of analyst consensus. Then, the function $g(\cdot)$ models the expected return only using the features of approximated consensus from the previous step, creating a “concept-bottleneck” within the prediction model. This empirical design is predicated on the understanding that both researchers in empirical asset pricing and financial analysts share the objective of assessing a firm’s value and discerning the factors that influence these valuations. While analysts often have access to broader datasets, including some predictive signals that may not be publicly available or included in this article, asset pricing panel data can represent a information subset in a decent quality by providing a comprehensive and quantifiable measures of firm’s fundamentals and macroeconomic conditions that are crucial for the approximation of the consensus, as shown in the empirical results later on.
@@
$\mathbb{E}_{t}\!\left[R_{i,t+h}\right]=g\!\left(\hat{C}_{i,t}\,;\,\theta\right).$ (3)
@@
Although it is designed as intended, given that neural networks are well-recognized as “universal approximators”, the model can allow any scenarios and consequences as outcomes, that don’t necessarily align with the economic theories. To overcome the limitation of the proposed prediction model, we apply stabilized optimization approaches proposed in the machine learning literature, such as regularization and scheduling. Such techniques are expected to function as “universal constraints”, achieving both practical performances and theoretical rigor. See Internet Appendix C.3 for detailed discussions and experimental settings.
@@ 2.2 Estimation
$\mathcal{L}=\mathcal{L}_{R}+\lambda\,\langle\textbf{1},\mathcal{L}_{C}\rangle,$ (4)
where the “return loss” $\mathcal{L}_{R}$ is formulated as,
$\mathcal{L}_{R}(\phi,\theta)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Bigl(R_{i,t+h}-g\!\left(f\!\left(\mathcal{I}^{f}_{i,t},\,\mathcal{I}^{m}_{t}\,;\,\phi\right)\,;\,\theta\right)\Bigr)^{2},$ (5)
and the “consensus loss” $\mathcal{L}_{C}$ is formulated as,
$\mathcal{L}_{C}(\phi)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Bigl(C_{i,t}-f\!\left(\mathcal{I}^{f}_{i,t},\,\mathcal{I}^{m}_{t}\,;\,\phi\right)\Bigr)^{2}.$ (6)
$\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ are cross-sectional mean squared errors (MSE) from a standard pooled OLS estimator, where $\lambda$ is a hyperparameter that assigns weight to the consensus loss, providing additional flexibility into the empirical design of the model.
@@
The strategy of jointly learning the consensus and excess return offers several advantages. Firstly, it tends to yield higher performance metrics due to the synergistic learning of interconnected variables. An alternative method might involve independent optimization, where the $f(\cdot)$ and $g(\cdot)$ are trained independently in two separate steps. However, this segmented approach often fails to capture the potential inter-dependencies between the consensus estimates and the resulting excess returns. Furthermore, since the information set $\mathcal{I}^{f}_{i,t}$ and $\mathcal{I}^{m}_{t}$ are not included in equation (5), the training of $g(\cdot)$ entirely depends on the quality of the extracted signals in approximated consensus, which makes training with the loss function $\mathcal{L}_{R}$ extremely challenging.
Secondly, it provides deeper insights and more intuitive understanding of the underlying financial dynamics. Independently learning $f(\cdot)$ using the equation (6) is not a novel concept and aligns with the existing literature supporting the evidence of the rational expectations hypothesis. As discussed in previous sections, the set of predictor signals used in this study is regarded to contain a significant amount of information sufficient to make “rational” expectations, [^4] Appendix D.5.2 formally evaluates this claim by examining the consensus-only specification corresponding to $\lambda\to\infty$ in Equation (4), showing that the model learns analysts’ consensus variables remarkably well (out-of-sample $R^{2}=30.30\%$) even without any return-prediction objective. This validates the architectural design of the consensus-bottleneck and provides empirical support for the rational expectations interpretation underlying the model. which simplifies the problem of approximating the opinions of individuals, compared to predicting future returns of assets. However, since analysts perform their analysis as their job, they must think and act beyond being merely “rational”; they must be “professional”. Therefore, we posit that professional and successful analysts strive to make their estimates predict not only “macroeconomic” consequences but also “firm-specific” outcomes. More specifically, proficient analysts will make decisions that better predict the future returns of a firm’s stocks.
@@ 2.3 Model architecture
In the proposed model, the consensus module is designed as an arbitrary feedforward network, while the prediction module is restricted to a simple linear regression that receives consensus variables as inputs and yields the expected excess return. This design choice is critical for enhancing interpretability in particular. When both modules are complex feedforward networks with multiple hidden layers, the advantage of using a consensus-based approach diminishes since it creates two separate black-box models from a single black-box model.
@@
$\mathrm{GELU}(x)=x\cdot\mathbb{P}(X\leq x)=x\cdot\Phi(x),$
where $\Phi(x)$ the cumulative distribution function for Gaussian distribution $X\sim N(0,\sigma^{2})$. GELU was first introduced by hendrycks2016gaussian as an alternative of Rectified linear units (ReLU) (nair2010rectified). Figure 2 shows that GELU permits some small interval for negative inputs to propagate through subsequent layers. See Internet Appendix C for a review of the literature on activation functions and justification for the selection of GELU.
@@
$H^{(1)}$ $=\mathrm{GELU}\!\left(W_{0}\!\left(\mathcal{I}^{f}\oplus\mathcal{I}^{m}\right)+b_{0}\right),$
$H^{(2)}$ $=\mathrm{GELU}\!\left(W_{1}H^{(1)}+b_{1}\right),$
@@
$g\!\left(f\!\left(\mathcal{I}^{f},\,\mathcal{I}^{m}\,;\,\phi\right)\,;\,\theta\right)=W^{r}_{n}\,f\!\left(\mathcal{I}^{f},\,\mathcal{I}^{m}\,;\,\phi\right)+b^{r}_{n}.$
@@ 3.1 Data description
In this section, we provide the brief explanations on the dataset and the sampling splitting scheme employed for empirical studies. The dataset comes from four distinct sources, which are all publicly available at the moment. Firstly, we obtain open-source asset pricing panel data from ChenZimmermann2021, available to download on their website (https://www.openassetpricing.com/). [^5] Data from ChenZimmermann2021 undergoes several preprocessing steps including lagging, data sampling, data imputation, and rank normalization, as detailed in Internet Appendix B. It comprises 114 firm-level predictors consisting of diverse financial metrics such as accounting figures, 13F filings, trading activities, and derivatives data.
@@
The final merged dataset consists of samples spanning from January 1994 to December 2023, with total 605,722 samples from 4,683 U.S. companies. Detailed descriptions of the dataset components and their respective sources are provided in Internet Appendix E.
@@ 3.2 Extracting macroeconomic state variables via autoencoder
To incorporate macroeconomic dynamics into the conditional expectation function $\mathbb{E}_{t}[R_{i,t+h}]$ defined in equation (2), we encode the aggregate information set $\mathcal{I}_{t}^{m}$ through an autoencoder-based representation. While macroeconomic variables are often dismissed in cross-sectional asset pricing due to their perceived homogeneity across firms, we argue that macro context exerts differentiated influence through sectoral dynamics, capital structure sensitivity, and behavioral channeling. However, the sheer volume and redundancy of macroeconomic indicators, particularly those sourced from databases such as FRED-MD, pose significant challenges for model training. Including hundreds of highly correlated variables not only increases the risk of overfitting but also dilutes the learning signal by overwhelming the model with noise and irrelevant information.
@@
Dimensionality reduction techniques have long been used in financial modeling to address such issues. Principal Component Analysis (PCA) has served as a standard tool for extracting latent factors from large panels of macroeconomic variables (ludvigson2007empirical), while extensions such as Sparse PCA and Independent Component Analysis (ICA) have been applied to improve factor interpretability and reduce multicollinearity (fan2016projected; erichson2020sparse). More recently, deep learning approaches—particularly autoencoders—have gained traction in the asset pricing literature for their ability to capture nonlinear interactions and extract economically meaningful latent structures from noisy, high-dimensional data (chen2024deep; gu2021autoencoder). These methods have proven effective in modeling complex macro-financial dynamics that traditional linear techniques may fail to uncover. [^7] Appendix D.4.2 demonstrates that replacing the autoencoder with a 32-factor PCA markedly weakens out-of-sample return predictability, despite both approaches delivering similar accuracy in reconstructing analysts’ consensus. This divergence highlights the advantage of nonlinear compression in capturing macroeconomic structure relevant for pricing.
@@
Formally, let the macroeconomic input at time $t$ be $\mathbf{x}_{t}\in\mathbb{R}^{D}$, where $D=123$. The encoder $\mathcal{E}_{\phi}(\cdot)$ maps this input to a latent representation $\mathbf{z}_{t}\in\mathbb{R}^{d}$: [^8] Empirically, setting the latent dimension to $d=32$ yields the best out-of-sample performance (see Internet Appendix D.4.1).
$\mathbf{z}_{t}=\mathcal{E}_{\phi}\!\left(\mathbf{x}_{t}\right).$
The decoder $\mathcal{D}_{\theta}(\cdot)$ reconstructs the input:
$\hat{\mathbf{x}}_{t}=\mathcal{D}_{\theta}\!\left(\mathbf{z}_{t}\right),$
@@
$\mathcal{L}_{\text{AE}}(\theta,\phi)=\frac{1}{T}\sum_{t=1}^{T}\left\|\mathbf{x}_{t}-\mathcal{D}_{\theta}\!\bigl(\mathcal{E}_{\phi}(\mathbf{x}_{t})\bigr)\right\|_{2}^{2}.$
@@
Formally, the latent representation $\mathbf{z}_{t}$ learned through the autoencoder serves as an empirical proxy for the macroeconomic information set $\mathcal{I}_{t}^{m}$ introduced in equation (2). In this context, $\mathbf{z}_{t}$ functions as a compressed, data-driven approximation of the macroeconomic state observable to investors at time $t$. This design allows the CB-APM to integrate the high-dimensional macroeconomic information set into a tractable latent representation, ensuring that firm-level forecasts remain conditioned on a parsimonious yet informative depiction of the aggregate economic environment. By mapping $\mathcal{I}_{t}^{m}$ into $\mathbf{z}_{t}$, the model effectively operationalizes the theoretical information set within a learnable structure, thereby linking the empirical implementation of the macro encoder to the conditional expectation framework defined in equation (2).
As illustrated in Figure 3, this framework visualizes the overall data pipeline of the CB-APM, depicting how macroeconomic inputs are encoded, compressed, and subsequently integrated with firm-level characteristics for return prediction. The figure serves as a conceptual representation clarifying the interaction between the macro autoencoder and the return-prediction module. The full architecture details of the autoencoder, including hidden-layer configurations and activation functions, are provided in Internet Appendix C.3. The empirical findings underscore the importance of representing macroeconomic regimes in shaping cross-sectional return dynamics and highlight the utility of neural representation learning in extracting economically salient signals from high-dimensional macro data. At each expanding-window step, the autoencoder is trained only on macro data available up to the window end date, and the encoder is then used to compute $\mathbf{z}_{t}$ for that window’s validation and test months, thereby preventing look-ahead bias.